### v1

# Self-Supervised Learning in Event Sequences: A Comparative Study and Hybrid Approach of Generative Modeling and Contrastive Learning

Abstract
This study investigates self-supervised learning techniques to obtain representations of Event Sequences (EvS). It is a key modality in various applications, including but not limited to banking, e-commerce, and healthcare.
We perform a comprehensive study of generative and contrastive approaches in self-supervised learning, applying them both independently. We find that there is no single supreme method. Consequently, we explore the potential benefits of combining these approaches. To achieve this goal, we introduce a novel method that aligns generative and contrastive embeddings as distinct modalities, drawing inspiration from contemporary multimodal research.
Generative and contrastive approaches are often treated as mutually exclusive, leaving a gap for their combined exploration. Our results demonstrate that this aligned model performs at least on par with, and mostly surpasses, existing methods and is more universal across a variety of tasks. Furthermore, we demonstrate that self-supervised methods consistently outperform the supervised approach on our datasets.
1 Introduction
Motivation. Event Sequences (EvS) are widely used in various real-world applications, including medicine Waring et al. (2020), biology Valeri et al. (2023), social medial analysis Liguori et al. (2023), fault diagnosis  Shao et al. (2019), churn prediction  Jain et al. (2021); Sudharsan and Ganesh (2022), customer segmentation  Carnein and Trautmann (2019), fraud detection  Xie et al. (2022) and more Zhuzhel et al. (2023); Fursov et al. (2021). As a result, there is a growing need to effectively model such data. Most self-supervised methods are focused on images, text, speech, or time-series. In Computer Vision (CV), contrastive learning methods have achieved state-of-the-art results as a pre-training strategy before supervised fine-tuning He et al. (2021); Chen et al. (2020); Caron et al. (2021). In Natural Language Processing (NLP), most modern models rely on generative pre-training Radford and Narasimhan (2018); Devlin et al. (2019). However, these approaches are largely unexplored for EvS.
The aim of our work is (1) to study generative and contrastive self-supervised approaches for pre-training and representation learning in the domain of EvS, (2) and to determine if they can complement each other. We study two ways to combine these approaches. For the first, we use a combined loss function. For the second, drawing inspiration from contemporary multimodal studies Radford et al. (2021); Zhai et al. (2023); Moskvoretskii and Kuznetsov (2023) we consider differently pre-trained models as distinct modalities and employ alignment techniques.
Contributions.
We are the first to our knowledge to examine the generative modeling for pre-training in the domain of EvS. We introduce a novel method called the Multimodal-Learning Event Model (MLEM) that aligns two pre-training strategies. Specifically, we use a pre-trained contrastive model to align a generative model. Our results have demonstrated that, on average, the aligned model outperforms any single method across a diverse range of tasks and datasets. Furthermore, our results indicate that self-supervised methods for pre-training outperform supervised approaches on all the datasets we have examined.
Furthermore, our study uncovers two significant insights. First, we find that the generative approach consistently achieves superior performance on tasks related to the Temporal Point Process (TPP), such as predicting the next event type and event time.
Second, we find that most of the methods are robust to perturbations along the time axis. However, they significantly degrade in performance on a synthetic dataset where the time component plays a crucial role. This observation suggests that some event sequences can be sufficiently classified using a ”Bag of Words” approach, where the order of events is not as important.
We provide the source code for all the experiments described in this paper. [^1] The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
2 Related work
In the following section, we will review various generative, contrastive, hybrid, and other related methods and works.
Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al. (2019); Raffel et al. (2023) and CV He et al. (2021); Assran et al. (2023); Bachmann et al. (2022). In NLP one of the generative pre-training procedures is the next token prediction used in BERT Devlin et al. (2019).
In Lin et al. (2022) authors study generative modeling of EvS for TPP tasks. Similarly to NLP, they use next-event prediction as their main training objective. While EvS has been studied in such a generative context, the authors in Lin et al. (2022) did not specifically investigate generative modeling as a representation of learning or pre-training. The similarities between autoregressive modeling in TPP and the success of generative approaches in NLP have prompted us to consider the potential for applying generative-style self-supervised learning and pre-training to the EvS domain.
Contrastive methods are a widely recognized in He et al. (2020); Grill et al. (2020). These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects. In CoLES Babaev et al. (2022) authors study contrastive approach for EvS. In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users. We utilize this method to examine the efficacy of the contrastive approach in our study.
Hybrid methods. Our work falls into the category of hybrid self-supervised approaches Qi et al. (2023); Yang et al. (2023); Lazarow et al. (2017); Oquab et al. (2023); Zhou et al. (2022); Kim et al. (2021); Kim and Ye (2022); Shukla and Marlin (2021). Few studies have merged generative and contrastive methods, typically focusing on loss combination. DINOv2 Oquab et al. (2023) updates the traditional contrastive loss from DINO Caron et al. (2021) with iBOT’s Zhou et al. (2022) reconstruction loss through Masked Image Modeling. Similarly, GCRL Kim et al. (2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive. EBCLR Kim and Ye (2022) also adopts this approach in the context of energy-based models, using a sum of contrastive and generative losses.
Other hybrid methods focus on combining generative and supervised learning Liu and Abbeel (2020). One such method, mTAND Shukla and Marlin (2021), has been applied for EvS classification.
Supervised learning is commonly used for the EvS classification task. Some works focus on modeling the underlying dynamics of the EvS process. For example, irregular samples can be modeled by employing dynamic modeling with Ordinary Differential Equations Rubanova et al. (2019). In contrast, SeqNAS Udovichenko et al. (2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process. A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate an RNN encoder in our model. Furthermore, the authors propose benchmark datasets for EvS classification. We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
Figure 1: A scheme for contrastive pre-training and representation learning is illustrated using two different sequences,{$S_{1},S_{2}$}, where $S\in\mathcal{R}^{l\times f}$ and $l$ is a sequence and sequence length, respectively. The sub-sequences {$S_{1}^{{}^{\prime}},S_{2}^{{}^{\prime}},S_{1}^{{}^{\prime\prime}},S_{2}^{{}^{\prime\prime}}$} are sampled using a subsequence sampler described in Section 4.1, and the latent representations {$h_{1}^{{}^{\prime}},h_{2}^{{}^{\prime}},h_{1}^{{}^{\prime\prime}},h_{2}^{{}^{\prime\prime}}$}, where $h\in\mathcal{R}^{m}$, are produced by the contrastive encoder model $\mathcal{E}_{c}$. In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$. The term $L^{con}$ denotes the contrastive loss.
Figure 2: In our generative method, the bottleneck encoder $\mathcal{E}_{g}$ encodes the entire sequence $S\in\mathcal{R}^{l\times f}$ into $h\in\mathcal{R}^{m}$, $l$ denotes sequence lengths and $f$ and $m$ corresponding hidden sizes. In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$. For generation, we employ a transformer decoder with recursive generation conditioned on $h$. The term $L^{LM}$ denotes the reconstruction loss, and $[BOS]$ represents the Beginning Of Sequence token.
Figure 3: MLEM approach for EvS. We illustrate it with two different sequences {$S_{1},S_{2}$}. Both the contrastive CON and generative GEN encoders map the sequences {$S_{1},S_{2}$} into corresponding {$h_{1}^{c},h_{2}^{c},h_{1}^{g},h_{2}^{g}$} latent representations. We align the representations from different models using SigLIP and compute the alignment loss, $L^{align}$ to train GEN. Simultaneously we train GEN encoder with another reconstruction objective $L^{LM}$. In the end, only the GEN encoder is used as the final model for fine-tuning and obtaining representations. Throughout the procedure, we use a pre-trained CON encoder with frozen weights.
3 Event Sequences Preliminaries
Datasets used in this work can be described as sets of sequences: $C=\{(S_{1},y_{1}),(S_{2},y_{2}),\ldots\}$, where $y_{i}$ is an attribute corresponding to the entire sequence or some target. Each $S_{i}=((t^{1}_{i},\Upsilon^{1}_{i}),(t^{2}_{i},\Upsilon^{2}_{i}),\ldots)$ is a sequence of events $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$, where $t_{i}^{j}$ represents the time when the event $x_{i}^{j}$ occurred and $t_{i}^{j}\leq t^{j+1}_{i}$. The set of sub-events $\Upsilon^{j}_{i}=\{k^{1},k^{2},\ldots\}$ describes the event $x_{i}^{j}$. It is important to note that in this work, the terms Event Sequences (EvS) and Irregularly Sampled Time-series (ISTS) are used interchangeably, as the occurrence of measurement or sampling for the latter can be seen as an event.
4 Methodology
4.1 Contrastive Learning
Contrastive Learning encodes a sequence $S$ into a compact representation $f_{e}:S_{i}\rightarrow h_{i}\in\mathcal{R}^{m}$, by bringing positive pairs (i.e., semantically similar objects) closer to each other in the embedding space, while pushing negative pairs (i.e., dissimilar objects) further apart. In the CoLES Babaev et al. (2022) framework, the authors suggest constructing a set of positive pairs by sampling sub-sequences from a sequence $S_{i}\rightarrow\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$, where each element in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ is shorter than $S_{i}$ and elements in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ may intersect, the number of sub-sequences is not limited to two. We adopt this approach and refer to it as the subsequence sampler in Figure 1. Further, we utilize loss function (1) from Hadsell et al. (2006). The overall procedure is illustrated in Figure 1.
$\begin{split}L^{con}&=\frac{1}{|C|}\sum_{i=1}^{|C|}\sum_{j=1}^{|C|}\Bigg{(}z_{ij}\frac{1}{2}||h_{i}-h_{j}||_{2}^{2}\\ &+(1-z_{ij})\frac{1}{2}\max\{0,\rho-||h_{i}-h_{j}||_{2}\}^{2}\Bigg{)}\end{split}$ (1)
Here, $z_{ij}\in\{0,1\}$ denotes if objects are semantically similar, $h$ is the sequence embedding, $\rho$ the minimal margin between dissimilar objects and $C=\{S_{1},S_{2},\ldots\}$ is a set of sequences.
4.2 Generative Modeling
We employ generative modeling to train our bottleneck encoder. Although our methodology is similar to Sequence-to-Sequence models, it differs as we use a bottleneck encoder to derive a single vector representation for the entire sequence.
For the encoder, we utilize RNN, which extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder Vaswani et al. (2017). The decoder reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism, in an auto-regressive manner. This entire process is illustrated in Figure 2.
To train the model we use the next event prediction objective, which is similar to training language models. Nonetheless, as each event in the sequence $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$ contain multiple sub-events $\Upsilon=\{k^{1},k^{2},\ldots\}$, we need to reconstruct all of them. To this end, we apply the cross-entropy loss for each categorical $k$ and the Mean Squared Error (MSE) loss for each real-valued $k$. The final loss is a cumulative sum of the losses for all elements in $\Upsilon=\{k^{1},k^{2},\ldots\}$, plus the MSE loss for time $t$. We denote this loss as $L^{LM}$.
Intuitively, this procedure requires our encoder to develop a representation informative enough for the decoder to accurately reconstruct the entire sequence. Since we map all our sequences $S$ into $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
All details related to the model can be found in Section 5.3.
Method ABank Age PhysioNet Pendulum TaoBao
ROC-AUC Accuracy ROC-AUC MSE ROC-AUC
Supervised $0.768\pm 0.000$ $0.602\pm 0.005$ $0.790\pm 0.021$ $\underline{0.33\pm 0.02}$ $0.684\pm 0.002$
Contrastive $0.729\pm 0.033$ $0.638\pm 0.007$ $\mathbf{0.815\pm 0.013}$ $\mathbf{0.26\pm 0.02}$ $0.679\pm 0.003$
Generative $\underline{0.788\pm 0.003}$ $0.639\pm 0.007$ $0.787\pm 0.007$ $\mathbf{0.26\pm 0.03}$ $\mathbf{0.695\pm 0.004}$
Naïve $0.658\pm 0.020$ $0.638\pm 0.007$ $0.759\pm 0.024$ $\mathbf{0.26\pm 0.04}$ $0.691\pm 0.002$
MLEM $\mathbf{0.790\pm 0.000}$ $\mathbf{0.642\pm 0.005}$ $0.780\pm 0.004$ $\mathbf{0.26\pm 0.05}$ $\mathbf{0.695\pm 0.002}$
Table 1: Evaluation of self-supervised methods fine-tuned using supervised loss. ABank, PhysioNet and TaoBao are reported with ROC-AUC, AGE reported with accuracy and Pendulum with MSE.
4.3 Aligning generative encoder
To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we treat them as distinct modalities and employ a contrastive aligning procedure inspired by CLIP Radford et al. (2021). We utilize the recent SigLIP loss Zhai et al. (2023) for this purpose. The resulting aligned encoder model is referred to as the Multimodal-Learning Event Model (MLEM).
Overall MLEM training procedure is exactly the same as for the generative model described in Section 4.2, except the alignment, which incurs an additional loss (2) resulting in the total loss (3). Training MLEM requires a pre-trained contrastive encoder. Using this encoder we train a generative model and align its embeddings with the embeddings produced by the contrastive encoder. Contrastive encoder weights are not updated during training. Both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding hidden states $h^{g}$ and $h^{c}$, then the goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away. This procedure is illustrated in Figure 3. Below is the alignment loss we use:
$L^{align}=\frac{1}{|C|}\sum_{i=1}^{|C|}\sum_{j=1}^{|C|}\underbrace{\log\frac{1}{1+e^{z_{ij}(-t\cdot{h^{g}_{i}}\cdot h^{c}_{j}+b)}}}_{L^{align}_{ij}}$ (2)
Here, $z_{ij}\in\{-1,1\}$ indicates if a pair of embeddings originated from the same sequence $S$, $t$ and $b$ denote the temperature and bias, respectively, both are learnable parameters.
Total loss to train MLEM:
$L=\alpha L^{LM}(S,\hat{S})+\beta L^{align}(\mathcal{E}_{g}(S),\mathcal{E}_{c}(S))$ (3)
Here, $S$ and $\hat{S}$ denote original and reconstructed sequences, $\alpha$ and $\beta$ are hyperparameters that adjust the strength of alignment.
All technical details can be found in Section 5.3.
4.4 Naïve method
As a baseline, we also examine a straightforward, or Naïve, approach that merges generative and contrastive methods. This is achieved by incorporating a contrastive objective into the generative model, akin to methodologies used in prior studies Kim et al. (2021); Oquab et al. (2023). The final loss is a weighted sum of objectives $L=\alpha L^{LM}+\beta L^{con}$. One of the downsides of this implementation is that we can not utilize full-length sequences because sub-sequence sampling is required for contrastive procedure. Further, we compare Naïve method with MLEM and show that while it is very close by design it leads to inferior results.
5 Experiments
5.1 Datasets
In our study, we use five EvS datasets formally outlined in Section 3.
  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions and TaoBao is user activity.
  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality.
  • Pendulum is a synthetic dataset simulating pendulum motion with the objective of predicting its length using a sequence of coordinates. This dataset was created specifically to evaluate a model’s capability to effectively incorporate the time component in sequence modeling.
Each dataset contains a range of features, both numerical and categorical. All datasets, with the exception of ABank, consist of irregularly sampled time steps. Comprehensive details for each dataset are available in the supplementary materials.
ABank Age PhysioNet Pendulum TaoBao
ROC-AUC TPP Acc. Acc. TPP Acc. ROC-AUC TPP MSE MSE TPP MSE ROC-AUC TPP Acc.
Linear Probing
Contrastive $0.678\pm 0.003$ $0.37$ $0.623\pm 0.010$ $0.28$ $\mathbf{0.819\pm 0.003}$ $1.05$ $\mathbf{0.37\pm 0.02}$ $0.8$ $0.680\pm 0.001$ $0.22$
Generative $0.753\pm 0.003$ $0.43$ $0.610\pm 0.004$ $0.3$ $0.759\pm 0.010$ $\mathbf{0.93}$ $0.74\pm 0.09$ $0.8$ $\mathbf{0.689\pm 0.005}$ $\mathbf{0.35}$
Naïve $0.703\pm 0.002$ $0.37$ $0.602\pm 0.003$ $\mathbf{0.31}$ $0.733\pm 0.005$ $\mathbf{0.93}$ $0.59\pm 0.07$ $0.8$ $0.680\pm 0.003$ $0.32$
MLEM $\mathbf{0.757\pm 0.000}$ $\mathbf{0.43}$ $\mathbf{0.627\pm 0.000}$ $0.29$ $0.762\pm 0.010$ $\mathbf{0.93}$ $0.41\pm 0.01$ $\mathbf{0.79}$ $0.683\pm 0.003$ $\mathbf{0.35}$
Non-linear Probing
Contrastive $0.691\pm 0.004$ $0.37$ $0.629\pm 0.010$ $0.28$ $\mathbf{0.822\pm 0.001}$ $1.05$ $\mathbf{0.37\pm 0.00}$ $0.8$ $0.686\pm 0.001$ $0.22$
Generative $0.758\pm 0.003$ $0.43$ $0.618\pm 0.004$ $0.3$ $0.764\pm 0.006$ $\mathbf{0.93}$ $0.63\pm 0.02$ $0.8$ $0.684\pm 0.002$ $\mathbf{0.35}$
Naïve $0.704\pm 0.005$ $0.37$ $0.608\pm 0.002$ $\mathbf{0.31}$ $0.774\pm 0.008$ $\mathbf{0.93}$ $0.57\pm 0.05$ $0.8$ $\mathbf{0.690\pm 0.002}$ $0.32$
MLEM $\mathbf{0.759\pm 0.003}$ $\mathbf{0.43}$ $\mathbf{0.634\pm 0.005}$ $0.29$ $0.780\pm 0.001$ $\mathbf{0.93}$ $0.40\pm 0.01$ $\mathbf{0.79}$ $0.688\pm 0.002$ $\mathbf{0.35}$
Table 2: Evaluation of self-supervised methods using linear and non-linear probing for downstream tasks and TPP tasks. The table reports the mean and standard deviation calculated from three different seeds for downstream task metrics. For TPP tasks, only the mean values from three seeds are presented, as the standard deviation was consistently $0.00$. In the table, the best-performing values are highlighted in bold, while the second-best values are underlined.
5.2 Evaluation objectives
To evaluate the effectiveness of self-supervised training strategies, our study focuses on two key aspects: the quality of the embeddings and the effectiveness of fine-tuning the entire network, following self-supervised studies Dubois et al. (2024, 2021).
5.2.1 Main Objectives
To assess the quality of embeddings, we primarily rely on metrics from downstream tasks. This involves utilizing both linear and non-linear probing methods. For non-linear probing, we employ Gradient Boosting algorithms with LGBM Ke et al. (2017) applied directly to the embeddings.
We evaluate embeddings on several tasks, including the prediction of an attribute or a target $y_{i}$ given the entire $S_{i}$ and the prediction of consequent event $x^{j}$ given a set of $S_{i}=\{x_{i}^{1},\ldots,x^{j-1}_{i}\}$. The second objective addresses the TPP task, which involves predicting the type of the next event or the time of the next event’s arrival. We have processed each dataset such that the target is either the category of the next event (for datasets that include this feature) or the time until the next event (for datasets without a primary event category).
5.2.2 Secondary Objectives
We incorporate additional metrics to assess the quality of embeddings. Specifically, we utilize anisotropy and intrinsic dimension, drawing on insights from previous research Nakada and Imaizumi (2020); Razzhigaev et al. (2023). In Section 6.3, we compare our findings with the results presented in the aforementioned works.
Anisotropy assesses the non-uniform distribution of embeddings in space, providing insights into the contextualization of our embedding. Lower anisotropy in embeddings has been linked to better model performance Ait-Saada and Nadif (2023). In line with the approach used in Razzhigaev et al. (2023), we compute anisotropy as the ratio of the highest singular value to the sum of all singular values:
$Anisotropy(H)=\frac{\sigma_{1}^{2}}{\sum\limits_{i}\sigma_{i}^{2}}$
The intrinsic dimension evaluates the optimal dimensionality of data, shedding light on the core information captured by the embeddings. To determine the intrinsic dimension, we employ the method proposed in Facco et al. (2017). This method examines how the volume of an $n$-dimensional sphere changes with dimension $d$. Further details are available in the original paper or in Razzhigaev et al. (2023).
5.3 Models
Feature embeddings. To transform features into a numerical vector, we use an embedding layer for categorical features and linear projection for numerical ones. For ABank, Age and TaoBao we set the dimension of each feature to 32. For Pendulum and PhysioNet dimension is set to 8. Additionally, we use the time difference between events as a separate feature.
For consistency in our experiments, we use the same encoder architecture across all training strategies and datasets. We use a single-layer GRU with a hidden size of 512, we take the last hidden state as sequence embedding. The GRU is selected for its proven effectiveness in encoding time-ordered sequences and its extensive application in relevant literature, providing a reliable baseline for our study Rubanova et al. (2019); Tonekaboni et al. (2021); Yoon et al. (2019); Udovichenko et al. (2024).
For Supervised model, we use the aforementioned encoder with a linear head.
Contrastive learning approach is described at 4.1 and uses the same encoder as mentioned above. Furthermore, we integrate a Feed Forward Projector atop the GRU for loss calculation, a technique commonly adopted in several studies for enhanced performance Grill et al. (2020); Oquab et al. (2023).
For the generative modeling, we employ a vanilla transformer decoder configured with LayerNorm, consisting of 3 layers, 2 heads, a hidden size of 128, and a Feed Forward hidden size of 256. To ensure training stability, we also apply LayerNorm to the sequence embedding produced by the encoder. Additionally, a projection layer is used atop the decoder to predict each feature.
5.4 Training details
To maintain consistency in our experiment, we trained all models using identical parameters. Models were trained for 100 epochs on datasets with fewer than 100K sequences and for 40 epochs on larger datasets. The learning rate was set at 1e-3, weight decay at 3e-3, and the batch size was fixed at 128. For the MLEM model, we set $\alpha$ at 1 and $\beta$ at 10. We averaged the results across three different seeds to ensure reliability and reproducibility.
6 Results
6.1 Self-supervised pre-training for fine-tuning
Here we evaluate the effectiveness of self-supervised pre-training strategies prior to supervised fine-tuning the entire model. For fine-tuning, we employed pre-trained encoders with randomly initialized linear layers and trained the model for 10 epochs. The results presented in Table 1 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, with the exception of the highly sparse PhysioNet dataset. Furthermore, it is noteworthy that supervised learning without self-supervised pre-training consistently produces inferior results.
6.2 Embeddings probing
We conducted a comprehensive analysis of the quality of embeddings using both linear and non-linear probing techniques, results are presented in Table 2. We evaluated performance on various tasks described in Section 5.2.1, including EvS classification and regression tasks for the entire sequence, as well as TPP tasks. Despite some variations in absolute values, the overall trends and rankings of methods remain consistent across both probing strategies.
Our assessment reveals that neither the contrastive nor the generative approach consistently outperforms the other. However, the MLEM model consistently demonstrates superior performance across most datasets. Notably, when MLEM is not the top performer, it often ranks second, highlighting its versatility. By effectively integrating the advantages of both contrastive and generative techniques, this approach consistently delivers strong overall performance across a range of datasets.
Figure 4: Mean percentage change averaged across datasets. The X-axis represents the dropout probability, while the Y-axis indicates the mean change in the linear probing downstream metric, compared to the metric with no dropout.
6.3 Anisotropy and Intrinsic Dimension
Method Anisotropy $\downarrow$ Intrinsic Dimension $\uparrow$
Contrastive $0.11\pm 0.04$ $10.15\pm 6.12$
Generative $0.08\pm 0.03$ $14.86\pm 10.12$
Naïve $0.07\pm 0.04$ $11.26\pm 6.51$
MLEM $\mathbf{0.06\pm 0.03}$ $\mathbf{15.86\pm 9.02}$
Table 3: Average anisotropy and intrinsic dimension across datasets
We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 5.2.2. As a result, we show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets. This finding may indicate that our pre-training strategy effectively enables the embeddings to encapsulate a greater amount of information and ensures their uniform distribution in space. The results are presented in Table 3.
Contrary to the findings in Dubois et al. (2024), we did not observe any correlation between intrinsic dimension and downstream performance metrics. Similarly, we did not find any correlation for anisotropy, which aligns with the aforementioned work.
Figure 5: Correlation plots for anisotropy and intrinsic dimension with normalized performance.
Figure 6: The t-SNE visualizations showcase embeddings resulting from various pre-training strategies. The top row displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset. Each point represents a sequence $S_{i}$ from a given dataset, colored accordingly to the corresponding attribute $y_{i}$. For Age, there are 4 classes and 2 classes for TaoBao
6.4 Visualization
The findings related to anisotropy and intrinsic dimension align with the t-SNE visualization of the Age and TaoBao datasets shown in Figure 6. While contrastive embeddings typically form distinct clusters, the generative approaches and MLEM provide a more complex embedding space.
7 Embedding Robustness
Our study also explores the robustness of model embeddings to specific types of perturbations. We focus on examining the impact of dropout and the shuffling of events within one sequence on the performance of the embeddings on downstream tasks.
7.1 Sensitivity to data omission
To investigate the effects of data omission, we applied a random dropout strategy, removing entire events from a sequence along the time axis, with varying probabilities $p_{dropout}\in\{0.1,0.3,0.5,0.7\}$. This allowed us to examine the impact of different levels of data removal on the embeddings, as seen in Figure 4, by measuring the average percentage decrease in downstream task metrics.
MLEM’s performance slightly decreases at $0.1$ dropout probability and more so at higher levels. This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on one specific sample. This property can be used in a sensitivity analysis or related applications.
Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline but actually surpassed their original performance at a dropout rate of 0.5. To draw reliable conclusions based on this finding, further work is required with a larger number of datasets and various models.
7.2 Sensitivity to perturbations
In subsequent tests, we examined the impact of perturbing events within a sequence on downstream task performance. The results, presented in Table 4, show that perturbing sequence elements had minimal to no effect on the results. This suggests that current models may be employing a ’Bag-of-Words’ approach and considering a set of events rather than the sequential nature of the data. One might assume that this discrepancy is due to not suitable models or training strategies. However, this hypothesis is not supported by a decline in performance on the pendulum dataset, where the time component is crucial and the same models were used.
This raises the question of whether we could further enhance model performance by explicitly considering the sequential nature of the data, or if this is solely a property of the datasets themselves.
Method ABank Age PhysioNet Pendulum TaoBao
Contrastive +0.47% -1.93% -0.37% -367.57% -0.15%
Generative -0.93% +0.16% -6.19% -91.89% -0.29%
Naive -1.42% +0.99% -3.00% -147.46% 0.00%
MLEM -1.45% -0.16% -6.56% -285.37% +0.29%
Table 4: Percentage change in linear probing performance after shuffling events order
8 Computational complexity
To obtain the MLEM encoder, we need to train two models, which is not desirable. However, to demonstrate the practicality of our method, we compare MLEM with SeqNas Udovichenko et al. (2024) in terms of performance and computational efficiency. SeqNas employs an architectural search to identify the most optimal supervised architecture for the task at hand. We posit that SeqNas represents an upper bound for downstream metrics. Table 5 presents results from the two largest datasets, examined by us, demonstrating that MLEM achieves performance near the upper bound set by SeqNas, but with significantly lower computational cost.
Method ABank Age
ROC-AUC GPU hours Accuracy GPU hours
MLEM $0.790\pm 0.000$ $25$ $0.642\pm 0.005$ $9$
SeqNas $0.7963\pm 0.001$ $288$ $0.645\pm 0.002$ $80$
Table 5: Comparison of downstream metrics and computational costs for training and fine-tuning MLEM versus SeqNas.
9 Discussion
We observe that in our settings, MLEM outperforms the Naïve method, despite their initial similarities. One possible explanation for the success of our method can be attributed to viewing MLEM as a Knowledge Distillation Hinton et al. (2015) procedure. However, further work is required to confirm this hypothesis.
We observed that data perturbation does not significantly impact model performance, suggesting a near-time-invariance. This could be attributed to the nature of the examined datasets or the absence of appropriate modeling approaches. Both hypotheses suggest the potential for future research in time-sensitive modeling.
We also observe consistent underperformance of methods that include generative approach on the PhysioNet dataset. This could be attributed to the high rate of missing values, which complicates the generation of accurate predictions for subsequent steps. Furthermore, the relatively small size of the dataset may be a contributing factor, given that generative modeling is generally dependent on large amounts of data.
10 Conclusion
We proposed a novel approach to combine multiple self-supervised strategies and demonstrated that this approach yields superior results compared to any single self-supervised method. We believe this opens up a new direction for enhancing self-supervised learning.
Our study revealed that self-supervised pre-trained models, especially those with generative pre-training, outperform purely supervised methods, with MLEM showing the most promise. This suggests potential benefits in exploring different pre-training and architectural options.
In our work, we are the first, to our knowledge, to apply generative modeling to event sequences for representation learning and pre-training. Our findings show that neither generative nor contrastive pre-training clearly surpasses the other in effectiveness. To address this, we developed the MLEM method, which innovatively merges contrastive learning with generative modeling. MLEM aligns these two approaches as separate modalities, inspired by recent advances in multimodal research, and it demonstrates enhanced performance on most datasets for both primary and secondary goals.

### v3

@@
-Self-Supervised Learning in Event Sequences: A Comparative Study and Hybrid Approach of Generative Modeling and Contrastive Learning
+MLEM: Generative and Contrastive Learning as Distinct Modalities for Event Sequences
 Abstract
-This study investigates self-supervised learning techniques to obtain representations of Event Sequences (EvS).
-It is a key modality in various applications, including but not limited to banking, e-commerce, and healthcare.
-We perform a comprehensive study of generative and contrastive approaches in self-supervised learning, applying them both independently.
-We find that there is no single supreme method.
-Consequently, we explore the potential benefits of combining these approaches.
-To achieve this goal, we introduce a novel method that aligns generative and contrastive embeddings as distinct modalities, drawing inspiration from contemporary multimodal research.
-Generative and contrastive approaches are often treated as mutually exclusive, leaving a gap for their combined exploration.
-Our results demonstrate that this aligned model performs at least on par with, and mostly surpasses, existing methods and is more universal across a variety of tasks.
-Furthermore, we demonstrate that self-supervised methods consistently outperform the supervised approach on our datasets.
+This study explores the application of self-supervised learning techniques for event sequences.
+It is a key modality in various applications such as banking, e-commerce, and healthcare.
+However, there is limited research on self-supervised learning for event sequences, and methods from other domains like images, texts, and speech may not easily transfer.
+To determine the most suitable approach, we conduct a detailed comparative analysis of previously identified best-performing methods.
+We find that neither the contrastive nor generative method is superior.
+Our assessment includes classifying event sequences, predicting the next event, and evaluating embedding quality.
+These results further highlight the potential benefits of combining both methods.
+Given the lack of research on hybrid models in this domain, we initially adapt the baseline model from another domain.
+However, upon observing its underperformance, we develop a novel method called the Multimodal-Learning Event Model (MLEM).
+MLEM treats contrastive learning and generative modeling as distinct yet complementary modalities, aligning their embeddings.
+The results of our study demonstrate that combining contrastive and generative approaches into one procedure with MLEM achieves superior performance across multiple metrics.
+**footnotetext: These authors contributed equally to this work
 1 Introduction
@@
 (2019).
-However, these approaches are largely unexplored for EvS.
-The aim of our work is (1) to study generative and contrastive self-supervised approaches for pre-training and representation learning in the domain of EvS, (2) and to determine if they can complement each other.
-We study two ways to combine these approaches.
-For the first, we use a combined loss function.
-For the second, drawing inspiration from contemporary multimodal studies Radford et al.
-(2021); Zhai et al.
-(2023); Moskvoretskii and Kuznetsov (2023) we consider differently pre-trained models as distinct modalities and employ alignment techniques.
-Contributions.
-We are the first to our knowledge to examine the generative modeling for pre-training in the domain of EvS.
-We introduce a novel method called the Multimodal-Learning Event Model (MLEM) that aligns two pre-training strategies.
-Specifically, we use a pre-trained contrastive model to align a generative model.
-Our results have demonstrated that, on average, the aligned model outperforms any single method across a diverse range of tasks and datasets.
-Furthermore, our results indicate that self-supervised methods for pre-training outperform supervised approaches on all the datasets we have examined.
-Furthermore, our study uncovers two significant insights.
-First, we find that the generative approach consistently achieves superior performance on tasks related to the Temporal Point Process (TPP), such as predicting the next event type and event time.
-Second, we find that most of the methods are robust to perturbations along the time axis.
-However, they significantly degrade in performance on a synthetic dataset where the time component plays a crucial role.
-This observation suggests that some event sequences can be sufficiently classified using a ”Bag of Words” approach, where the order of events is not as important.
-We provide the source code for all the experiments described in this paper. [^1] The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
-2 Related work
-In the following section, we will review various generative, contrastive, hybrid, and other related methods and works.
-Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al.
-(2019); Raffel et al.
-(2023) and CV He et al.
-(2021); Assran et al.
-(2023); Bachmann et al.
+Self-Supervised Learning (approaches have not been extensively explored for EvS.
+However, applying existing SSL methods directly to EvS may be unreasonable or even impossible due to several key differences between EvS and other well-studied domains: a.
+Multiple Features, events in EvS can be described by multiple features, which adds complexity to the learning process.
+b.
+Categorical and Numerical Values, event features in EvS may include both categorical and numerical values, further complicating modeling.
+c.
+Non-Uniform Temporal Spacing, events in EvS are not uniformly spaced in time, which requires models to handle irregular temporal patterns.
+d.
+Temporal Proximity, it is to assumed that events occurring in close temporal proximity share relevant information.
+Given these unique characteristics of EvS, further research is needed to explore SSL approaches specifically tailored to ES and address these challenges.
+We describe a formal structure of EvS in Section 2.
+The aim of our work is to answer the following questions: Q1: Which self-supervised approach is superior for ES modeling: generative or contrastive?
+Q2: Can we efficiently combine generative and contrastive approaches into one procedure to obtain more robust results?
+The Contributions.
+To answer these questions, we (1) study established generative, contrastive, and hybrid approaches.
+(2) To the best of our knowledge, we are the first to examine generative pre-training, for EvS classification, and hybrid methods in the domain of EvS in general.
+(3) We introduce a new technique named MLEM, inspired by contemporary multimodal studies, such as (Radford et al., 2021; Zhai et al., 2023; Moskvoretskii and Kuznetsov, 2023).
+MLEM allows for the combination of various SSL approaches by treating differently pre-trained models as distinct modalities.
+More precisely, MLEM aligns the contrastive and generative embeddings.
+We demonstrate that such combination achieves superior performance on average across multiple datasets and metrics compared to existing methods.
+Evaluation employs a diverse set of datasets to cover various aspects of the problem space: three real-world transactional datasets from the SeqNAS benchmark (Udovichenko et al., 2024), the widely-used medical dataset Physionet 2012, and a simulated dataset of pendulum motion designed to examine explicit time dependencies.
+This selection ensures that our findings are robust across different types of data and tasks.
+To comprehensively measure performance we employ a set of metrics, including downstream tasks of regression or classification of sequences, prediction of the next event type or timing, intrinsic dimension, and anisotropy.
+Baselines.
+To ensure a clear and fair comparison, we have selected only the top-performing model from each setting for our analysis, comparison of baseline approaches is presented in Table 3 and Table 3.
+These include models from recent studies on contrastive learning (Babaev et al., 2022), generative modeling in EvS (Lin et al., 2022), and hybrid approaches from other domains Kim et al.
 (2021).
-In NLP one of the generative pre-training procedures is the next token prediction used in BERT Devlin et al.
-(2019).
-In Lin et al.
-(2022) authors study generative modeling of EvS for TPP tasks.
-Similarly to NLP, they use next-event prediction as their main training objective.
-While EvS has been studied in such a generative context, the authors in Lin et al.
-(2022) did not specifically investigate generative modeling as a representation of learning or pre-training.
-The similarities between autoregressive modeling in TPP and the success of generative approaches in NLP have prompted us to consider the potential for applying generative-style self-supervised learning and pre-training to the EvS domain.
-Contrastive methods are a widely recognized in He et al.
-(2020); Grill et al.
-(2020).
-These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects.
-In CoLES Babaev et al.
-(2022) authors study contrastive approach for EvS.
-In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users.
-We utilize this method to examine the efficacy of the contrastive approach in our study.
-Hybrid methods.
-Our work falls into the category of hybrid self-supervised approaches Qi et al.
-(2023); Yang et al.
-(2023); Lazarow et al.
-(2017); Oquab et al.
-(2023); Zhou et al.
-(2022); Kim et al.
-(2021); Kim and Ye (2022); Shukla and Marlin (2021).
-Few studies have merged generative and contrastive methods, typically focusing on loss combination.
-DINOv2 Oquab et al.
-(2023) updates the traditional contrastive loss from DINO Caron et al.
-(2021) with iBOT’s Zhou et al.
-(2022) reconstruction loss through Masked Image Modeling.
-Similarly, GCRL Kim et al.
-(2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive.
-EBCLR Kim and Ye (2022) also adopts this approach in the context of energy-based models, using a sum of contrastive and generative losses.
-Other hybrid methods focus on combining generative and supervised learning Liu and Abbeel (2020). One such method, mTAND Shukla and Marlin (2021), has been applied for EvS classification.
-Supervised learning is commonly used for the EvS classification task.
-Some works focus on modeling the underlying dynamics of the EvS process.
-For example, irregular samples can be modeled by employing dynamic modeling with Ordinary Differential Equations Rubanova et al.
-(2019).
-In contrast, SeqNAS Udovichenko et al.
-(2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process.
-A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate an RNN encoder in our model.
-Furthermore, the authors propose benchmark datasets for EvS classification.
-We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
-Figure 1: A scheme for contrastive pre-training and representation learning is illustrated using two different sequences,{$S_{1},S_{2}$}, where $S\in\mathcal{R}^{l\times f}$ and $l$ is a sequence and sequence length, respectively.
-The sub-sequences {$S_{1}^{{}^{\prime}},S_{2}^{{}^{\prime}},S_{1}^{{}^{\prime\prime}},S_{2}^{{}^{\prime\prime}}$} are sampled using a subsequence sampler described in Section 4.1, and the latent representations {$h_{1}^{{}^{\prime}},h_{2}^{{}^{\prime}},h_{1}^{{}^{\prime\prime}},h_{2}^{{}^{\prime\prime}}$}, where $h\in\mathcal{R}^{m}$, are produced by the contrastive encoder model $\mathcal{E}_{c}$.
-In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$.
-The term $L^{con}$ denotes the contrastive loss.
-Figure 2: In our generative method, the bottleneck encoder $\mathcal{E}_{g}$ encodes the entire sequence $S\in\mathcal{R}^{l\times f}$ into $h\in\mathcal{R}^{m}$, $l$ denotes sequence lengths and $f$ and $m$ corresponding hidden sizes.
-In our work, we utilize the GRU (Gated Recurrent Unit) and extract the last hidden state as $h$.
-For generation, we employ a transformer decoder with recursive generation conditioned on $h$.
-The term $L^{LM}$ denotes the reconstruction loss, and $[BOS]$ represents the Beginning Of Sequence token.
-Figure 3: MLEM approach for EvS.
-We illustrate it with two different sequences {$S_{1},S_{2}$}.
-Both the contrastive CON and generative GEN encoders map the sequences {$S_{1},S_{2}$} into corresponding {$h_{1}^{c},h_{2}^{c},h_{1}^{g},h_{2}^{g}$} latent representations.
-We align the representations from different models using SigLIP and compute the alignment loss, $L^{align}$ to train GEN.
-Simultaneously we train GEN encoder with another reconstruction objective $L^{LM}$.
-In the end, only the GEN encoder is used as the final model for fine-tuning and obtaining representations.
-Throughout the procedure, we use a pre-trained CON encoder with frozen weights.
+We focus on these models because they are clearly superior in their categories, making the inclusion of lesser models unnecessary.
+Figure 1: An example of one sequence $S_{i}$ with some categorical and real-valued sub-events $\{k^{1},k^{2},\ldots\}$.
+We provide the source code for all the experiments described in this paper.***The source code of our work is publicly available at https://github.com/VityaVitalich/MLEM.
 2 Event Sequences Preliminaries
@@
 It is important to note that in this work, the terms Event Sequences (EvS) and Irregularly Sampled Time-series (ISTS) are used interchangeably, as the occurrence of measurement or sampling for the latter can be seen as an event.
+Figure 1 depicts an example of one sequence.
+Figure 2: Schematic Models Overview: A) The Contrastive Model divides sequence $S_{i}$ into subsequences $\{S_{i}^{\prime},S_{i}^{\prime\prime}\}$, encodes them into latent vectors $\{h_{i}^{\prime},h_{i}^{\prime\prime}\}$, and applies contrastive loss $L_{\text{con}}$.
+B) The Generative Model autoregressively reconstructs sequence $S$ cross-attended to latent vector $h$, derived from $S$ via a bottleneck encoder.
+Language model loss $L_{\text{LM}}$ is applied to reconstructed sequence $\hat{S}$.
+C) The MLEM Hybrid Model integrates generative and contrastive approaches, treating them as distinct modalities.
+It generates latent vectors $h_{i}^{g}$ and $h_{i}^{c}$ from a learned generative encoder and frozen, pretrained contrastive encoder respectively.
+An alignment loss $L_{\text{align}}$ is utilized to align vectors from the same sequence while separating those from different sequences.
+Language modeling loss $L_{\text{LM}}$ is applied as in generative model.
 3 Methodology
 3.1 Contrastive Learning
-Contrastive Learning encodes a sequence $S$ into a compact representation $f_{e}:S_{i}\rightarrow h_{i}\in\mathcal{R}^{m}$, by bringing positive pairs (i.e., semantically similar objects) closer to each other in the embedding space, while pushing negative pairs (i.e., dissimilar objects) further apart.
-In the CoLES Babaev et al.
-(2022) framework, the authors suggest constructing a set of positive pairs by sampling sub-sequences from a sequence $S_{i}\rightarrow\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$, where each element in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ is shorter than $S_{i}$ and elements in $\{S_{i}^{{}^{\prime}},S_{i}^{{}^{\prime\prime}}\}$ may intersect, the number of sub-sequences is not limited to two.
-We adopt this approach and refer to it as the subsequence sampler in Figure 1.
-Further, we utilize loss function (1) from Hadsell et al.
-(2006).
-The overall procedure is illustrated in Figure 1.
-$\begin{split}L^{con}&=\frac{1}{|C|}\sum_{i=1}^{|C|}\sum_{j=1}^{|C|}\Bigg{(}z_{ij}\frac{1}{2}||h_{i}-h_{j}||_{2}^{2}\\ &+(1-z_{ij})\frac{1}{2}\max\{0,\rho-||h_{i}-h_{j}||_{2}\}^{2}\Bigg{)}\end{split}$ (1)
-Here, $z_{ij}\in\{0,1\}$ denotes if objects are semantically similar, $h$ is the sequence embedding, $\rho$ the minimal margin between dissimilar objects and $C=\{S_{1},S_{2},\ldots\}$ is a set of sequences.
+For our contrastive learning approach, we employ COLES Babaev et al.
+(2022) framework, as it has shown superior performance over other contrastive methods.
+It learns to encode the sequence $S$ into a latent vector $h\in\mathcal{R}^{m}$ by bringing sub-sequences of the same sequence closer in the embedding space, while pushing sub-sequences from different sequences further apart.
+The overall procedure is illustrated in Figure 2 and formal definition is given in Appendix Contrastive learning
 3.2 Generative Modeling
-We employ generative modeling to train our bottleneck encoder.
-Although our methodology is similar to Sequence-to-Sequence models, it differs as we use a bottleneck encoder to derive a single vector representation for the entire sequence.
-For the encoder, we utilize RNN, which extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder Vaswani et al.
-(2017).
-The decoder reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism, in an auto-regressive manner.
+For our generative model, we consider the best performing setting from GNTPP (Lin et al., 2022) and adapt it to our task.
+For the encoder, we utilize RNN, as it shows dominance to our task in recent studies (Udovichenko et al., 2024).
+It extracts the final hidden state $h$, which is subsequently passed into a Transformer decoder (Vaswani et al., 2017).
+The decoder autoregressively reconstructs the entire sequence, conditioned on $h$ via the Cross-Attention mechanism.
 This entire process is illustrated in Figure 2.
-To train the model we use the next event prediction objective, which is similar to training language models.
+To train the model, we use the next event prediction objective, which is similar to training language models.
 Nonetheless, as each event in the sequence $x^{j}_{i}=(t^{j}_{i},\Upsilon^{j}_{i})$ contain multiple sub-events $\Upsilon=\{k^{1},k^{2},\ldots\}$, we need to reconstruct all of them.
@@
 Intuitively, this procedure requires our encoder to develop a representation informative enough for the decoder to accurately reconstruct the entire sequence.
-Since we map all our sequences $S$ into $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
-All details related to the model can be found in Section 5.3.
+Since we map all our sequences $S$ into, $h\in\mathcal{R}^{m}$ we call all encoders bottleneck encoders.
+Table 1: EvS modeling and choice of baselines
+Supervised Contrastive Generative
+Paper SeqNAS CoLES Generative TPP
+2024 2022 2022
+Best base models GRU GRU REV-ATT, LSTM
+Table 2: Hybrid SSL approaches
+Sum of Losses (Naïve) Multimodal
+Paper DINO 2023,
+iBOT 2022 MLEM
+GCRL 2021
+Domain Images and Text EvS
+Table 3: Evaluation of methods fine-tuned with supervised loss for downstream task.
 Method ABank Age PhysioNet Pendulum TaoBao
@@
 MLEM $\mathbf{0.790\pm 0.000}$ $\mathbf{0.642\pm 0.005}$ $0.780\pm 0.004$ $\mathbf{0.26\pm 0.05}$ $\mathbf{0.695\pm 0.002}$
-Table 1: Evaluation of self-supervised methods fine-tuned using supervised loss. ABank, PhysioNet and TaoBao are reported with ROC-AUC, AGE reported with accuracy and Pendulum with MSE.
-4.3 Aligning generative encoder
-To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we treat them as distinct modalities and employ a contrastive aligning procedure inspired by CLIP Radford et al.
+3.3 Naïve Hybrid Method
+We consider weighted summation of contrastive and generative loss as baseline hybrid model, as it showed dominant performance in previous studies (Kim et al., 2021).
+We call this approach Naïve since it is a straightforward combination of losses $L=\alpha L^{LM}+\beta L^{con}$.
+3.4 MLEM
+In our method, named MLEM, we propose to combine contrastive learning and generative modeling by treating them as distinct modalities with further alignment.
+To align the embeddings obtained through generative modeling with those acquired through contrastive learning, we employ a contrastive aligning procedure inspired by CLIP Radford et al.
 (2021).
@@
 (2023) for this purpose.
-The resulting aligned encoder model is referred to as the Multimodal-Learning Event Model (MLEM).
-Overall MLEM training procedure is exactly the same as for the generative model described in Section 4.2, except the alignment, which incurs an additional loss (2) resulting in the total loss (3).
-Training MLEM requires a pre-trained contrastive encoder.
-Using this encoder we train a generative model and align its embeddings with the embeddings produced by the contrastive encoder.
-Contrastive encoder weights are not updated during training.
-Both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding hidden states $h^{g}$ and $h^{c}$, then the goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away.
+Overall MLEM training procedure is nearly the same as for the generative model described in Section 3.2, except the alignment, which incurs an additional loss (1) resulting in the total loss (2).
+Unlike CLIP and other multimodal models, our approach involves pre-training only one modality.
+Specifically, we pre-train the contrastive encoder first and then train the generative encoder, aligning it with the fixed contrastive encoder.
+Formally, both generative $\mathcal{E}_{g}$ and contrastive $\mathcal{E}_{c}$ encoders receive a set of sequences $C=\{S_{1},S_{2},\ldots\}$ and map each sequence $S$ into corresponding latent vectors $h^{g}$ and $h^{c}$.
+The goal of MLEM alignment is to pull embeddings obtained from the same sequence closer to each other and to push embeddings from different sequences further away.
 This procedure is illustrated in Figure 2.
@@
 Here, $S$ and $\hat{S}$ denote original and reconstructed sequences, $\alpha$ and $\beta$ are hyperparameters that adjust the strength of alignment.
-All technical details can be found in Section 5.3.
-4.4 Naïve method
-As a baseline, we also examine a straightforward, or Naïve, approach that merges generative and contrastive methods.
-This is achieved by incorporating a contrastive objective into the generative model, akin to methodologies used in prior studies Kim et al.
-(2021); Oquab et al.
-(2023).
-The final loss is a weighted sum of objectives $L=\alpha L^{LM}+\beta L^{con}$.
-One of the downsides of this implementation is that we can not utilize full-length sequences because sub-sequence sampling is required for contrastive procedure.
-Further, we compare Naïve method with MLEM and show that while it is very close by design it leads to inferior results.
 4 Experiments
 4.1 Datasets
-In our study, we use five EvS datasets formally outlined in Section 3.
-  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions and TaoBao is user activity.
-  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality.
+In our study, we use five established EvS datasets formally outlined in Section 2.
+  • A subset of EvS datasets from SeqNAS benchmark Udovichenko et al. (2024), more precisely: ABank, Age, TaoBao. ABank and Age are bank transactions, and TaoBao is user activity.
+  • PhysioNet 2012 comprises highly sparse medical event data, with the goal of predicting mortality Silva et al. (2012).
   • Pendulum is a synthetic dataset simulating pendulum motion with the objective of predicting its length using a sequence of coordinates.
 This dataset was created specifically to evaluate a model’s capability to effectively incorporate the time component in sequence modeling.
-Each dataset contains a range of features, both numerical and categorical.
-All datasets, with the exception of ABank, consist of irregularly sampled time steps.
-Comprehensive details for each dataset are available in the supplementary materials.
+The detailed description is provided in Appendix Detailed Dataset Information.
+Table 4: Evaluation of self-supervised methods using linear and non-linear probing for downstream and TPP tasks.
+The best-performing values are highlighted in bold, while the second-best values are underlined.
 ABank Age PhysioNet Pendulum TaoBao
@@
 MLEM $\mathbf{0.759\pm 0.003}$ $\mathbf{0.43}$ $\mathbf{0.634\pm 0.005}$ $0.29$ $0.780\pm 0.001$ $\mathbf{0.93}$ $0.40\pm 0.01$ $\mathbf{0.79}$ $0.688\pm 0.002$ $\mathbf{0.35}$
-Table 2: Evaluation of self-supervised methods using linear and non-linear probing for downstream tasks and TPP tasks.
-The table reports the mean and standard deviation calculated from three different seeds for downstream task metrics.
-For TPP tasks, only the mean values from three seeds are presented, as the standard deviation was consistently $0.00$.
-In the table, the best-performing values are highlighted in bold, while the second-best values are underlined.
 4.2 Evaluation objectives
-To evaluate the effectiveness of self-supervised training strategies, our study focuses on two key aspects: the quality of the embeddings and the effectiveness of fine-tuning the entire network, following self-supervised studies Dubois et al.
-(2024, 2021).
-5.2.1 Main Objectives
-To assess the quality of embeddings, we primarily rely on metrics from downstream tasks.
-This involves utilizing both linear and non-linear probing methods.
-For non-linear probing, we employ Gradient Boosting algorithms with LGBM Ke et al.
-(2017) applied directly to the embeddings.
-We evaluate embeddings on several tasks, including the prediction of an attribute or a target $y_{i}$ given the entire $S_{i}$ and the prediction of consequent event $x^{j}$ given a set of $S_{i}=\{x_{i}^{1},\ldots,x^{j-1}_{i}\}$.
-The second objective addresses the TPP task, which involves predicting the type of the next event or the time of the next event’s arrival.
-We have processed each dataset such that the target is either the category of the next event (for datasets that include this feature) or the time until the next event (for datasets without a primary event category).
-5.2.2 Secondary Objectives
-We incorporate additional metrics to assess the quality of embeddings.
-Specifically, we utilize anisotropy and intrinsic dimension, drawing on insights from previous research Nakada and Imaizumi (2020); Razzhigaev et al.
-(2023).
-In Section 6.3, we compare our findings with the results presented in the aforementioned works.
-Anisotropy assesses the non-uniform distribution of embeddings in space, providing insights into the contextualization of our embedding.
+To evaluate the effectiveness of self-supervised training strategies, we follow previous studies Dubois et al. (2024, 2021) and focus on two key aspects:
+  • Fine-Tuning Performance. The downstream metrics after fine-tuning the entire network.
+  • Embeddings Probing. The quality of the embeddings for downstream and TPP tasks tested via linear and non-linear probing.
+As well, we include additional objectives for embedding quality, that were discovered in previous studies Razzhigaev et al. (2023); Nakada and Imaizumi (2020) , that are connected with performance:
+  • Anisotropy.
+It assesses the non-uniformity of embeddings in space, providing insights into the contextualization of our embedding.
 Lower anisotropy in embeddings has been linked to better model performance Ait-Saada and Nadif (2023).
-In line with the approach used in Razzhigaev et al.
-(2023), we compute anisotropy as the ratio of the highest singular value to the sum of all singular values:
-$Anisotropy(H)=\frac{\sigma_{1}^{2}}{\sum\limits_{i}\sigma_{i}^{2}}$
-The intrinsic dimension evaluates the optimal dimensionality of data, shedding light on the core information captured by the embeddings.
-To determine the intrinsic dimension, we employ the method proposed in Facco et al.
-(2017).
-This method examines how the volume of an $n$-dimensional sphere changes with dimension $d$.
-Further details are available in the original paper or in Razzhigaev et al.
-(2023).
+  • Intrinsic dimension. It evaluates the optimal dimensionality of data, showing information captured by the embeddings.
+Objectives details are described in appendix Objectives description.
 4.3 Models
-Feature embeddings.
+Base model and Feature embeddings.
 To transform features into a numerical vector, we use an embedding layer for categorical features and linear projection for numerical ones.
@@
 Additionally, we use the time difference between events as a separate feature.
-For consistency in our experiments, we use the same encoder architecture across all training strategies and datasets.
-We use a single-layer GRU with a hidden size of 512, we take the last hidden state as sequence embedding.
-The GRU is selected for its proven effectiveness in encoding time-ordered sequences and its extensive application in relevant literature, providing a reliable baseline for our study Rubanova et al.
+We use GRU as an encoder because it has been proven effective in encoding time-ordered sequences  Rubanova et al.
 (2019); Tonekaboni et al.
@@
 (2024).
+Udovichenko et al.
+performs a Neural Architecture Search on various models and highlights that the presence of RNN blocks in the architecture is associated with higher performance on average.
+Lin et al.
+illustrate that the performance gain brought by history encoders is very small for TPP modeling.
+Based on these empirical observations, we believe that our design choice is reasonable.
+We it with a single-layer and a hidden size of 512, we take the last hidden state as sequence embedding.
+For consistency in our experiments, we use the same hyperparameters across all training strategies and datasets.
 For Supervised model, we use the aforementioned encoder with a linear head.
-Contrastive learning approach is described at 4.1 and uses the same encoder as mentioned above.
+Contrastive learning approach is described at 3.1 and uses the same encoder as mentioned above.
 Furthermore, we integrate a Feed Forward Projector atop the GRU for loss calculation, a technique commonly adopted in several studies for enhanced performance Grill et al.
@@
 The learning rate was set at 1e-3, weight decay at 3e-3, and the batch size was fixed at 128.
-For the MLEM model, we set $\alpha$ at 1 and $\beta$ at 10.
+For the Hybrid models, we set $\alpha$ at 1 and $\beta$ at 10.
 We averaged the results across three different seeds to ensure reliability and reproducibility.
+The standard deviation could be not reported if negligible.
 5 Results
-6.1 Self-supervised pre-training for fine-tuning
-Here we evaluate the effectiveness of self-supervised pre-training strategies prior to supervised fine-tuning the entire model.
-For fine-tuning, we employed pre-trained encoders with randomly initialized linear layers and trained the model for 10 epochs.
-The results presented in Table 1 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, with the exception of the highly sparse PhysioNet dataset.
-Furthermore, it is noteworthy that supervised learning without self-supervised pre-training consistently produces inferior results.
+5.1 Fine-tuning performance
+First, we evaluate the effectiveness of fine-tuning the entire model after pre-training procedure.
+The results presented in Table 3 indicate that, among all the evaluated pre-training methods, MLEM consistently achieves the most favorable results across all datasets, except for the highly sparse PhysioNet dataset.
+Furthermore, it is noteworthy that solely supervised learning consistently produces inferior results.
 5.2 Embeddings probing
 We conducted a comprehensive analysis of the quality of embeddings using both linear and non-linear probing techniques, results are presented in Table 4.
-We evaluated performance on various tasks described in Section 5.2.1, including EvS classification and regression tasks for the entire sequence, as well as TPP tasks.
-Despite some variations in absolute values, the overall trends and rankings of methods remain consistent across both probing strategies.
-Our assessment reveals that neither the contrastive nor the generative approach consistently outperforms the other.
-However, the MLEM model consistently demonstrates superior performance across most datasets.
-Notably, when MLEM is not the top performer, it often ranks second, highlighting its versatility.
-By effectively integrating the advantages of both contrastive and generative techniques, this approach consistently delivers strong overall performance across a range of datasets.
-Figure 4: Mean percentage change averaged across datasets.
-The X-axis represents the dropout probability, while the Y-axis indicates the mean change in the linear probing downstream metric, compared to the metric with no dropout.
-6.3 Anisotropy and Intrinsic Dimension
+Our evaluation reveals that neither the contrastive nor the generative approach consistently outperforms the other.
+However, by effectively integrating the advantages of both contrastive and generative techniques, MLEM model demonstrates superior performance across most datasets on average.
+Notably, when MLEM is not the top performer, it is often ranked second, highlighting its versatility.
+Table 5: Average anisotropy and intrinsic dimension across datasets
 Method Anisotropy $\downarrow$ Intrinsic Dimension $\uparrow$
@@
 MLEM $\mathbf{0.06\pm 0.03}$ $\mathbf{15.86\pm 9.02}$
-Table 3: Average anisotropy and intrinsic dimension across datasets
-We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 5.2.2.
-As a result, we show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets.
+Table 6: Performance change in linear probing performance after shuffling events order, higher is better
+Method ABank Age PhysioNet Pendulum TaoBao
+Contrastive +0.47% -1.93% -0.37% -367.57% -0.15%
+Generative -0.93% +0.16% -6.19% -91.89% -0.29%
+Naive -1.42% +0.99% -3.00% -147.46% 0.00%
+MLEM -1.45% -0.16% -6.56% -285.37% +0.29%
+5.3 Anisotropy and Intrinsic Dimension
+We conducted an evaluation of anisotropy and the intrinsic dimensions of embeddings from various pre-training strategies, as detailed in 4.2.
+Results presented in Table 6 show that MLEM has the highest intrinsic dimension and the lowest anisotropy, on average, across all datasets.
 This finding may indicate that our pre-training strategy effectively enables the embeddings to encapsulate a greater amount of information and ensures their uniform distribution in space.
-The results are presented in Table 3.
-Contrary to the findings in Dubois et al.
-(2024), we did not observe any correlation between intrinsic dimension and downstream performance metrics.
-Similarly, we did not find any correlation for anisotropy, which aligns with the aforementioned work.
-Figure 5: Correlation plots for anisotropy and intrinsic dimension with normalized performance.
+However, we found no correlation between intrinsic dimension or anisotropy and downstream performance in our domain, as illustrated in correlation plot 4(b).
+While our findings align with the most recent study by Dubois et al.
+(2024) regarding anisotropy, they differ in that they observed a correlation between intrinsic dimension and performance.
 Figure 3: The t-SNE visualizations showcase embeddings resulting from various pre-training strategies.
-The top row displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset.
+The left portion displays the embeddings for the Age dataset, while the bottom row illustrates those for the TaoBao dataset.
 Each point represents a sequence $S_{i}$ from a given dataset, colored accordingly to the corresponding attribute $y_{i}$.
@@
 Our study also explores the robustness of model embeddings to specific types of perturbations.
-We focus on examining the impact of dropout and the shuffling of events within one sequence on the performance of the embeddings on downstream tasks.
+We focus on examining the impact of dropout and the shuffling of events within one sequence on the linear probing performance.
+We assess robustness with classification or regression tasks.
+In each scenario, we assess either the class of the whole sequence or the real value describing the whole sequence (pendulum length).
+We refer to this assessment as global and assume that this attribute does not change over the entire sequence (e.g., user age group).
+Attributes that change over time are referred to as local (e.g., amount of payment or next purchase category).
+6.1 Sensitivity to permutation
+We randomly permuted order of events along time axis and analyzed performance of pre-trained models.
+Models were trained without permutation.
+The results, presented in Table 6, show that the sequence order has only a small effect on most of the results.
+This suggests that current models may be employing a ’Bag-of-Words’ approach, treating the events as a set rather than considering the sequential nature of the data when determining certain global attributes.
+Additionally, we observed that the dataset’s degradation is greater when it has fewer categorical features.
+Despite the small degradation, it is still present, highlighting the need for modeling temporal relations.
+Furthermore, all the models demonstrate the ability to encode sequential data, as evidenced by a decline in performance on the pendulum dataset where the time component is crucial.
+The assessment of the effect of order permutation on the ability to model local attributes is similar to the TPP task, and we leave it for the future work.
 6.2 Sensitivity to data omission
 To investigate the effects of data omission, we applied a random dropout strategy, removing entire events from a sequence along the time axis, with varying probabilities $p_{dropout}\in\{0.1,0.3,0.5,0.7\}$.
-This allowed us to examine the impact of different levels of data removal on the embeddings, as seen in Figure 4, by measuring the average percentage decrease in downstream task metrics.
-MLEM’s performance slightly decreases at $0.1$ dropout probability and more so at higher levels.
-This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on one specific sample.
+We report the average percentage decrease in linear probing performance, depicted in Figure 4(a).
+We notice that MLEM performance slightly decreases at $0.1$ dropout probability and more so at higher levels.
+This indicates that MLEM is more sensitive to dropout and requires the presence of all the samples rather than depending on a couple specific samples.
 This property can be used in a sensitivity analysis or related applications.
-Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline but actually surpassed their original performance at a dropout rate of 0.5.
-To draw reliable conclusions based on this finding, further work is required with a larger number of datasets and various models.
-7.2 Sensitivity to perturbations
-In subsequent tests, we examined the impact of perturbing events within a sequence on downstream task performance.
-The results, presented in Table 4, show that perturbing sequence elements had minimal to no effect on the results.
-This suggests that current models may be employing a ’Bag-of-Words’ approach and considering a set of events rather than the sequential nature of the data.
-One might assume that this discrepancy is due to not suitable models or training strategies.
-However, this hypothesis is not supported by a decline in performance on the pendulum dataset, where the time component is crucial and the same models were used.
-This raises the question of whether we could further enhance model performance by explicitly considering the sequential nature of the data, or if this is solely a property of the datasets themselves.
-Method ABank Age PhysioNet Pendulum TaoBao
-Contrastive +0.47% -1.93% -0.37% -367.57% -0.15%
-Generative -0.93% +0.16% -6.19% -91.89% -0.29%
-Naive -1.42% +0.99% -3.00% -147.46% 0.00%
-MLEM -1.45% -0.16% -6.56% -285.37% +0.29%
-Table 4: Percentage change in linear probing performance after shuffling events order
+Additionally, the generative modeling embeddings exhibited an intriguing behavior: they not only resisted the decline, but actually surpassed their original performance at a dropout rate of 0.5.
+We believe that such effect show high sensitivity to the presence of noisy or out-of-distribution events.
+Nevertheless, to draw reliable conclusions, further work is required.
+(a) Mean percentage change averaged across datasets.
+The X-axis represents the dropout probability, the Y-axis indicates the mean change in the linear probing, compared to the metric with no dropout.
+Error bars reflect 1 sigma deviation.
+(b) Correlation plots for anisotropy and intrinsic dimension with normalized performance.
 7 Computational complexity
@@
 (2024) in terms of performance and computational efficiency.
+Table 7: Comparison of downstream metrics and computational costs for training MLEM versus SeqNas.
+Method ABank Age
+ROC-AUC GPU hours Accuracy GPU hours
+MLEM $0.790\pm 0.000$ $47$ $0.642\pm 0.005$ $9$
+SeqNas $0.7963\pm 0.001$ $288$ $0.645\pm 0.002$ $80$
 SeqNas employs an architectural search to identify the most optimal supervised architecture for the task at hand.
@@
 Table 9 presents results from the two largest datasets, examined by us, demonstrating that MLEM achieves performance near the upper bound set by SeqNas, but with significantly lower computational cost.
-Method ABank Age
-ROC-AUC GPU hours Accuracy GPU hours
-MLEM $0.790\pm 0.000$ $25$ $0.642\pm 0.005$ $9$
-SeqNas $0.7963\pm 0.001$ $288$ $0.645\pm 0.002$ $80$
-Table 5: Comparison of downstream metrics and computational costs for training and fine-tuning MLEM versus SeqNas.
-9 Discussion
-We observe that in our settings, MLEM outperforms the Naïve method, despite their initial similarities.
-One possible explanation for the success of our method can be attributed to viewing MLEM as a Knowledge Distillation Hinton et al.
-(2015) procedure.
-However, further work is required to confirm this hypothesis.
-We observed that data perturbation does not significantly impact model performance, suggesting a near-time-invariance.
-This could be attributed to the nature of the examined datasets or the absence of appropriate modeling approaches.
-Both hypotheses suggest the potential for future research in time-sensitive modeling.
-We also observe consistent underperformance of methods that include generative approach on the PhysioNet dataset.
+Detailed compute budget reported in table 9
+8 Related work
+In the following section, we will review various generative, contrastive, hybrid, and other related methods and works.
+Generative methods for pre-training and representation learning have seen notable development in recent years in both NLP Radford and Narasimhan (2018); Devlin et al.
+(2019); Raffel et al.
+(2023) and CV He et al.
+(2021); Assran et al.
+(2023); Bachmann et al.
+(2022).
+In the EvS domain, the sole study by Lin et al.
+(2022) investigates various generative models for TPP tasks.
+Based on their findings, we have considered their best performing setting to serve as a generative model for our study, with slight modifications to fit our tasks.
+Contrastive methods are a widely recognized in He et al.
+(2020); Grill et al.
+(2020).
+These methods typically draw closer the embeddings of variations of the same object and distance them from those of different objects.
+In CoLES Babaev et al.
+(2022) authors study contrastive learning for EvS.
+In their work, the positive samples consist of subsequences derived from a single user’s events, while the negative samples are obtained from events of different users.
+We utilize this method as best performing along contrastive approaches.
+Hybrid methods.
+Our work falls into the category of hybrid self-supervised approaches Qi et al.
+(2023); Yang et al.
+(2023); Lazarow et al.
+(2017) Few studies have merged generative and contrastive methods, typically focusing on loss combination.
+DINOv2 Oquab et al.
+(2023) updates the traditional contrastive loss from DINO Caron et al.
+(2021) with iBOT’s Zhou et al.
+(2022) reconstruction loss through Masked Image Modeling.
+Similarly, GCRL Kim et al.
+(2021) combines these losses via a weighted sum, significantly outperforming models that are purely generative or contrastive.
+We follow the same methodology to form the hybrid baseline method with simply summing contrastive and generative losses.
+Supervised learning is commonly used for the EvS classification task.
+The most comprehensive recent paper SeqNAS Udovichenko et al.
+(2024) adopts Neural Architecture Search to identify optimal architectures without strong prior assumptions about the underlying process.
+A key finding is the effectiveness of Recurrent Neural Networks (RNNs) in modeling EvS, leading us to incorporate RNN encoder in our model.
+Furthermore, the authors propose benchmark datasets for EvS classification.
+We selected the largest datasets from the benchmark, as self-supervised methods require substantial data volumes.
+9 Limitations
+Physionet underperformane.
+We observe consistent underperformance of methods that include generative approach on the PhysioNet dataset.
 This could be attributed to the high rate of missing values, which complicates the generation of accurate predictions for subsequent steps.
 Furthermore, the relatively small size of the dataset may be a contributing factor, given that generative modeling is generally dependent on large amounts of data.
+Compute Resources.
+One limitation of our model is the requirement for a higher computational budget compared to previous methods.
+However, in the domain of event sequences, the primary challenge is typically achieving the highest metric rather than optimizing for computational efficiency.
+Datasets and Models.
+We selected the most relevant datasets to cover a broad spectrum of tasks and chose the best-performing models from previous studies to represent SSL methods.
+Nonetheless, the comparison could be improved by incorporating a wider variety of datasets and model architectures.
+Scalability.
+Although we aimed to ensure a fair comparison, the results may vary with different scales of data or models.
+However, observed trend show that more data leads to better results on MLEM and generative method.
 10 Conclusion
-We proposed a novel approach to combine multiple self-supervised strategies and demonstrated that this approach yields superior results compared to any single self-supervised method.
-We believe this opens up a new direction for enhancing self-supervised learning.
-Our study revealed that self-supervised pre-trained models, especially those with generative pre-training, outperform purely supervised methods, with MLEM showing the most promise.
-This suggests potential benefits in exploring different pre-training and architectural options.
-In our work, we are the first, to our knowledge, to apply generative modeling to event sequences for representation learning and pre-training.
-Our findings show that neither generative nor contrastive pre-training clearly surpasses the other in effectiveness.
-To address this, we developed the MLEM method, which innovatively merges contrastive learning with generative modeling.
-MLEM aligns these two approaches as separate modalities, inspired by recent advances in multimodal research, and it demonstrates enhanced performance on most datasets for both primary and secondary goals.
+To our knowledge, our study is the first to apply generative modeling to event sequences for representation learning and pre-training.
+We found that neither generative nor contrastive pre-training definitively outperforms the other.
+In response, we developed the MLEM method, which creatively combines contrastive learning with generative modeling.
+Inspired by recent advances in multimodal research, MLEM treats these approaches as separate modalities and consistently outperforms existing models on most datasets across both primary and secondary objectives.
+This finding suggests a promising new direction for enhancing self-supervised learning across various domains.
+Additionally, our research provides several key insights.
+Firstly, the best current self-supervised models do not utilize the sequential nature of data effectively.
+We also discovered that random event deletion does not adversely affect performance; in some cases, it actually enhances it.
+Secondly, unlike findings in other domains (Dubois et al., 2024; Ait-Saada and Nadif, 2023), we observed no correlation between anisotropy or intrinsic dimension and downstream performance.